# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 8 Noisy Group Testing
Up until now, we have assumed that the tests are noiseless, i.e. the test is positive iff there is at least one infection in the group that is tested. In this section, we extend our results to the case of noisy tests. We consider the symmetric noisy model introduced in Section 2.1, where each test result is flipped with a small probability $20<\delta<1/$ 2, independent of the other tests.
A simple and standard argument shows that if we repeat each test for $T_{R}>\frac{\log(\mu+H(X))}{1-4\delta(1-\delta)}$ times, and then take the majority, with high probability, we get the correct result for all the tests. So we incur an additional factor of $\frac{\log(\mu+H(X))}{1-4\delta(1-\delta)}$.
Another way to solve this issue is to calculate the posterior probability considering the fact that tests are noisy. The following theorem shows that with small modifications in Algorithm 1, we can find the infected set better than just repeating each test.
@@
Suppose the result of each test is flipped with probability $\delta<1/2$, $\delta<\epsilon$, and $c>1-\frac{1}{(1-\delta)^{1-\delta}\delta^{\delta}}$. There is an algorithm that performs $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log(1-c)-H(\delta)}H(X)+\frac{\log\mu}{1-4\delta(1-\delta)}\frac{\mu}{2\epsilon(1-2c)}$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-2\epsilon$.
@@
We first describe how we update the posterior when the tests are noisy. Consider an edge $e$ and a test $T\subseteq V$, where $e\cap T=\emptyset$. If $T$ is negative, there are two possibilities. Either it is not flipped, which happens with probability $1-\delta$ and $p(e)$ is updated to $\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, where $E^{\prime}$ is the set of edges that have no intersection with $T$, or the test is flipped and $p(e)$ is updated to 00. So the posterior probability for $e$ is $p(e)=(1-\delta)\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$ Similarly, if the test is positive, if it is not flipped $p(e)$ is updated to 00 and if it is flipped, $p(e)$ is updated to $\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, hence the posterior probability is $\delta\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$. With the same argument, we can obtain the posterior probability when $e\cap T\neq\emptyset$, and only the posterior for positive and negative tests are swapped.
Again by Markov Bound, with probability $1-\epsilon$ the number of infections is less than $\mu/\epsilon$. From now on, we allow for $\epsilon$ error, as $(1-\epsilon)^{2}>1-2\epsilon$, which is our target error. Now, instead of line 1 where the algorithm returns the single edge, if there is an edge with probability more than $1-\epsilon$, the algorithm returns it. Suppose $e^{*}$ is the target edge and $T$ tests are performed in the first phase. Then with high probability and by the new posterior obtained in the last part, we have $q_{e^{*}}=(1-\delta)^{(1-\delta)T}\delta^{\delta T}(\frac{1}{1-c})^{T}=((1-\delta)^{1-\delta}\delta^{\delta}\frac{1}{1-c})^{T}>(c^{\prime})^{T}$ where $c^{\prime}>1$. Hence it takes $\log{1/p(e)}/\log c^{\prime}$ tests for $q_{e^{*}}$ to become greater than $1-\epsilon$. Similar to Corollary 4.6, the number of tests in this phase is $\frac{1}{\log c^{\prime}}H(X)$
Now for the second phase, same as the previous part, we repeat each single test in line 1 and take the majority. Again, if we repeat each test $O(\frac{\log\mu}{1-4\delta(1-\delta)})$ times, with high probability we recover all the nodes correctly and hence we perform $\frac{\log\mu}{1-4\delta(1-\delta)}\frac{\mu}{2\epsilon(1-2c)}$ tests. ∎
Theorem 8.1 shows that our framework can handle the noise and update the posterior accordingly. The following theorem shows under the same definition of noise, we can achieve a result in the SNAGT setting by just repeating the tests.
@@
Suppose the result of each test is flipped with probability $\delta<1/2$, and for a graph $G$, the probability distribution $\mathcal{D}$ is such that $\Pr_{e\sim\mathcal{D}}(|\left|e\right|-\mu|>\epsilon\mu)=o(1)$ for any constant $\epsilon>0$. Then, there is an algorithm in the SNAGT setting that performs $O(\frac{\log[\mu\cdot H(X)+\mu\log n)]}{1-4\delta(1-\delta)}[\mu\cdot H(X)+\mu\log n)]$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
Proof.
We employ the same algorithm described in Theorem 7.1, but modify it to address noise by incorporating test repetition and majority voting. Specifically, each test is repeated $\frac{\log[\mu\cdot H(X)+\mu\log(n)]}{1-4\delta(1-\delta)}$ times, and the outcome of the test is determined by the majority vote among these repetitions.
By choosing the number of repetitions as above, we ensure that the probability of a single test producing an incorrect majority outcome is at most $\mathrm{err}\ll\frac{1}{\mu\cdot H(X)+\mu\log(n)}.$ Applying a union bound over all tests, the probability that any test is incorrect tends to zero as $n$ grows.
Consequently, with high probability, we may assume that all tests are effectively noiseless. Under this assumption, the situation reduces exactly to the noise-free setting of Theorem 7.1. Hence, the same performance guarantees and conclusions from that theorem apply, completing the proof. ∎

### v2

@@ 8 Noisy Group Testing
Up until now, we have assumed that the tests are noiseless, i.e. the test is positive iff there is at least one infection in the group that is tested. In this section, we extend our results to the case of noisy tests. We consider the symmetric noisy model introduced in Section 2.1, where each test result is flipped with a probability $20<\delta<1/$ 2, independent of the other tests.
If we repeat each test for at least $\Theta(\frac{\log n}{(1-2\delta)^{2}})$ times, and then take the majority verdict, using a standard application of Chernoff bound, one can show that we get the correct result for all the tests, with a probability that goes to $1$ as $n$ goes to infinity. So we incur an additional multiplicative factor of $\Theta(\frac{\log n}{(1-2\delta)^{2}})$ in the expected number of tests, and the upper bound on the expected number of tests for Algorithm 1 with each test repeated as above becomes $\Theta(\frac{\log n}{(1-2\delta)^{2}})O(H(X)+u)$ (Theorem 4.11 when $Pr[|e^{*}|>u]$ is $O(1/n)$). We show that Algorithm 1 can be modified to improve the upper bound on the expected number of tests by updating the posterior probability after each test to incorporate the impact of the noise in the tests and repeating the tests only in some steps:
@@
Suppose that the result of each test is flipped with probability $\delta<1/2$, and $21-{(1-\delta)^{1-\delta}\delta^{\delta}}<c<1/$ 2 . Also, $Pr[|e^{*}|>u]=O(1/n)$. Then, there is an algorithm that performs $\frac{1}{\log[1/(1-c)]-H(\delta)}H(X)+O(\frac{\log(g(n)u)}{(1-2\delta)^{2}}\frac{u}{(1-2c)})+\Theta(\log^{2}n/(1-2\delta)^{2})$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-O(1/\min(n,g(n))$. Here, $g(n)$ is any function that goes to infinity as $n\rightarrow\infty$.
Remark 8.2.
We first compare the guarantee on the number of tests in Theorem 8.1 with when noise is handled by repeating every test. Consider the case that $\delta$ does not increase with $n$ and $u$ is $o(n).$ Then $\delta$ can be replaced with its initial value in the upper bounds, which is a constant. Then the additional factor $\frac{1}{\log 1/(1-c)-H(\delta)}$ for the first term in the upper bound on the number of tests in Theorem 8.1, which is a constant and is therefore $o(\log n).$ The additional coefficient for $u$ in Theorem 8.1 is $\frac{\log[g(n)u]}{(1-2\delta)^{2}}$, which is therefore $o(\log n)$ since $g(n)$ can be any function that increases to infinity as $n$ increases to infinity. So the coefficient of both $H(X)$ and $u$ are asymptotically substantially smaller than the multiplicative factor for the case that tests are merely repeated. But the upper bound in Theorem 8.1 has an additional additive term of $\Theta(\log^{2}n)$ as compared to the bound when the tests are merely repeated. However, the additive term is not an impediment if $H(X)$ is $\Omega(\log n).$ To see this, first consider that $H(X)$ is both $\omega(\log n)$ and $O(\log^{2}n).$ The upper bound on the expected number of tests in Theorem 8.1 is $\Theta(\log^{2}n)$, whereas that when tests are merely repeated is $\omega(\log^{2}n).$ Now, consider that $H(X)$ is $\Omega(\log^{2}n)$. The upper bound on the expected number of tests in Theorem 8.1 is $\Theta(H(X)$, whereas that when tests are merely repeated is $\Theta(\log n)\Theta(H(X)).$ Now, consider the case that $H(X)$ is $\Theta(\log n).$ Then, the upper bounds in Theorem 8.1 and when we merely repeat tests are both $\Theta(\log^{2}n)$. The additional additive term in the upper bound in Theorem 8.1 is significant when $H(X)$ is $o(\log n)$, as by repeating each test $\Theta((\log n)/(1-2\delta)^{2})$ times, both terms in the total number of tests $\Theta(\frac{\log n}{(1-2\delta)^{2}})O(H(X)+u)$ are $o(\log^{2}n)$ Thus, the upper bound in Theorem 8.1 is $\Theta(\log^{2}n)$ whereas that when we only repeat each test is $o(\log^{2}n)$. In this regime, simple well-known testing algorithms can provide asymptotically vanishing error probability using limited number of tests in the basic noiseless case  [^4] For example, when the entropy is a constant, there are only a few number of edges with significant probability. Then the complement of each such edge can be tested in decreasing order of size of edges. For the first test that is negative, all edges that are not fully contained in the edge whose complement has been tested can be removed. The same process can be repeated for edges that are fully contained in this edge. The largest remaining edge at the end of this process is the target edge. The number of tests equals the number of edges identified initially as those with significant probability of being a target edge., thus even the basic scenario in this case may not be of significant interest any longer.
The coefficient for $H(X)$ in Theorem 8.1 is $\frac{1}{\log 1/(1-c)-H(\delta)}$. It is known that $\frac{1}{1-H(\delta)}H(X)$ is a lower bound on the number of tests and is an optimal multiplicative factor in independent group testing with symmetric noise. Thus, the $\log(1/(1-c))$ is appearing due to the strategies for accommodating correlations in states of nodes.
The theorem is not contingent upon $\delta$ being a constant, and allows $\delta$ to depend on $n$. In practice, for large communities (large $n$), the error probability can be lower as the local government (eg in big cities as opposed to rural counties) often has more resources and can therefore acquire more reliable test kits. The upper bound on the expected number of tests in Theorem 8.1 increases with increase in the probability of error $\delta$ between 00 and $1/2$ Specifically, when $\delta\rightarrow 1/2$, ${(1-\delta)^{1-\delta}\delta^{\delta}}\rightarrow 1/2.$ Thus, since $21-{(1-\delta)^{1-\delta}\delta^{\delta}}<c<1/$ 2, $c\rightarrow 1/2$ too. Then the coefficient of both $H(X)$, which is $\frac{1}{\log 1/(1-c)-H(\delta)}$, and the coefficient of $u$, which is $\frac{\log[g(n)u]}{(1-2\delta)^{2}}$ go to infinity. Thus, there is a tradeoff between error rate and guarantee on the number of tests.
We now prove Theorem 8.1.
@@
We modify Algorithm 1 as follows: 1) each test in line 1 and line 1 is repeated for $\Theta(\log n/(1-2\delta)^{2})$ and $\Theta(\frac{\log g[(n)u]}{(1-2\delta)^{2}})$ times respectively. Each test is deemed positive if the majority is positive, it is negative otherwise (if a test is not repeated the majority is the outcome in the lone instance). The tests in other lines in the algorithm (ie, line 1) are however not repeated as above. The posterior probability of each edge is updated based on the majority outcome of each test to incorporate the impact of the probabilistic flips in results. We also never perform more than $n$ tests, and the algorithm halts if and when that threshold is breached, it can halt even earlier due to the embedded decisions.
Updating the posterior:
We now describe how we update the posterior for a test. Consider an edge $e$ and a test $T\subseteq V$. Let $T^{o},T^{r}$ be the deemed and real results of the test respectively, the outcome of the test reveals $T^{o}$, but not $T^{r}.$ Let $T^{o}=a$, $q(e)=Pr(e=e^{*}\mid T^{o}=a)$, where $a\in\{-1,+1\}.$ Thus,
$q(e)=Pr(e=e^{*}\mid T^{o}=a,T^{r}=a)Pr(T^{r}=a\mid T^{o}=a)+Pr(e=e^{*}\mid T^{o}=a,T^{r}=b)Pr(T^{r}=b\mid T^{o}=a).$
But $[T^{r}=a\mid T^{o}=a]$ is equivalent to result not being flipped which has probability $1-\delta$, and $[T^{r}=b\mid T^{o}=a]$ is equivalent to the result being flipped, which has probability $\delta$. Moreover, $Pr(e=e^{*}\mid T^{o},T^{r})=Pr(e=e^{*}\mid T^{r})$. Thus, if $T^{o}=a$,
$q(e)=(1-\delta)Pr(e=e^{*}\mid T^{r}=a)+\delta Pr(e=e^{*}\mid T^{r}=b).$
Note that $p(e)=Pr(e=e^{*}\mid T^{r}=-1)*Pr(T^{r}=-1)+Pr(e=e^{*}\mid T^{r}=+1)*Pr(T^{r}=+1).$ Clearly then $\max(Pr(e=e^{*}\mid T^{r}=-1),Pr(e=e^{*}\mid T^{r}=+1))>0$, as otherwise $p(e)=0.$ Since $0<\delta<1,$ it follows that $q(e)>0,$ if $p(e)>0.$ Thus, if the posterior for an edge is positive to start with, it remains positive throughout. So no edge is ever eliminated, and the posterior for every edge always remains below $1$ (otherwise the posterior for other edges would reduce to 00). Thus, line 1 is the only place that the algorithm actually outputs the infected nodes.
Upper bounding the number of tests in line 1:
We first provide a lower bound on the posterior of the target edge, and then using the lower bound, we provide an upper bound on the number of tests in line 1.
Lower bounding the posterior: First let $e\cap T=\emptyset$, $Pr(e=e^{*}\mid T^{r}=-1)=\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, where $E^{\prime}$ is the set of edges that have no intersection with $T$, and $Pr(e=e^{*}\mid T^{r}=+1)=0.$ Thus, if $T^{o}$ is negative, $q(e)=(1-\delta)\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$. Note that $\sum_{f\in E^{\prime}}p_{f}\leq 1-c$: since $\sum_{f\in E^{\prime}}p_{f}=w(V\setminus T)$, and by the algorithm $T$ is tested before line 1 if $c\leq w(V\setminus T)\leq 1-c$. Thus, if $T^{o}$ is negative, $q(e)\geq(1-\delta)\frac{p(e)}{1-c}.$
Now let $e\cap T\neq\emptyset$. Then, $Pr(e=e^{*}\mid T^{r}=-1)=0$, and $Pr(e=e^{*}\mid T^{r}=+1)=\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, where $E^{\prime}$ is the set of edges that have a nonempty intersection with $T$. Now, by the algorithm, $c\leq w(V\setminus T)\leq 1-c$. Also, $w(E^{\prime})=1-w(V\setminus T)$, since $E^{\prime}$ consists of all edges which have at least one node in $T$ and therefore all edges that are not contained in $V\setminus T.$ Thus, since $w(V\setminus T)\geq c$, $w(E^{\prime})\leq 1-c$. Thus, if $T^{o}$ is negative, $q(e)=\delta\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}\geq\delta\frac{p(e)}{1-c}$.
Proceeding similarly, if $T^{o}$ is positive, 1) $q(e)\geq\delta\frac{p(e)}{1-c}$ if $e\cap T=\emptyset$, and 2) $q(e)\geq(1-\delta)\frac{p(e)}{1-c}$ if $e\cap T\neq\emptyset.$
Since $1-\delta\geq\delta,$ $q(e)\geq\delta\frac{p(e)}{1-c}$, in all instances.
Upper bound: Suppose $e^{*}$ is the target edge and $M$ tests are performed in line 1. Consider the tests for which 1) $e^{*}\cap T=\emptyset$ and $T^{o}$ is negative, or 2) $e^{*}\cap T\neq\emptyset$ and $T^{o}$ is positive. In each such test, since every node in $e^{*}$ is positive and no node outside $e^{*}$ is positive, the test result has not been flipped. In each such test $q(e^{*})\geq(1-\delta)\frac{p(e^{*})}{1-c}$ by the analysis in the previous paragraphs. Let the number of tests that have not been flipped be $M_{1}.$ It follows that after the $M$ tests, $q_{e^{*}}\geq p(e^{*})(1-\delta)^{M_{1}}\delta^{|M|-M_{1}}(\frac{1}{1-c})^{M}.$
Since the flips in results are i.i.d., by the Chernoff bound, we have $Pr[M_{1}<(1-\Delta)\nu]\leq e^{-\frac{\nu\Delta^{2}}{2}}$ where $\nu=\mathbf{E}[M]=(1-\delta)M$. Setting $\Delta=2\sqrt{\frac{\log n}{\nu}}$, we get
$Pr[M_{1}<\nu-2\sqrt{\log(n)\nu}]<e^{2\log n}=1/n^{2}.$
Hence, with probability more than $1-1/n^{2}$, we have $M_{1}\geq(1-\delta)M-\partial{M_{1}}$ where $\partial{M_{1}}=2\sqrt{\log(n)(1-\delta)M}$. Let the number of tests $M$ be greater than $\log^{2}n.$ Then, $\frac{\partial M_{1}}{M}\rightarrow 0$ as $n\rightarrow\infty$, thus $\partial M_{1}$ is $o(M)$ and $\frac{\partial M_{1}}{M}$ is also $O(1/\sqrt{\log n}).$
By the new posterior update obtained, we can now lower bound the posterior after $M$ tests with high probability. With probability $1-1/n^{2}$, we have
$q_{e^{*}}$ $\geq p(e^{*})(1-\delta)^{(1-\delta)M-\partial M}\delta^{\delta M+\partial M}(\frac{1}{1-c})^{M}$
$=p(e^{*})((\frac{\delta}{1-\delta}^{\frac{\partial M_{1}}{M}}(1-\delta)^{1-\delta}\delta^{\delta}\frac{1}{1-c})^{M}$
$>p(e^{*})(\frac{\delta}{1-\delta}^{\frac{\partial M_{1}}{M}}c^{\prime})^{M},$
where $c^{\prime}=(1-\delta)^{1-\delta}\delta^{\delta}\frac{1}{1-c}>1$, and $\lim_{n\rightarrow\infty}\frac{\delta}{1-\delta}^{\frac{\partial M_{1}}{M}}=1$. Note that no edge ever gets a probability of $1$, but after at most $\frac{\log(1/p(e^{*}))}{\log(c^{\prime})}+\log^{2}n$ tests, $q(e^{*})>1$ with probability at least $1-(1/n^{2}).$ Thus, the algorithm must have fewer than $\log^{2}n+\frac{\log(1/p(e^{*}))}{\log(c^{\prime})}$ tests in line 1 with probability at least $1-(1/n^{2})$; with the remaining probability it must have $n$ or fewer tests. Thus, given that $e^{*}$ is the target edge the expected number of tests in line 1 is at most $\log^{2}n+\frac{\log(1/p(e^{*}))}{\log(c^{\prime})}+(1/n).$
Since each edge $e$ is the target edge with probability $p(e)$, the expected number of tests $M$ in line 1 is at most $\frac{1}{\log c^{\prime}}H(X)+\log^{2}n+(1/n).$
In the above analysis we have implicitly assumed that the algorithm conducts tests in line 1 consecutively, but in reality it may move to line 1 from line 1, and return to line 1 every time the majority outcome in line 1 is positive. But we show later that whenever such return happens, the posterior for the target edge is multiplied by a factor of at least $(1-\delta)/(1-c)$ with a high probability, the probability that this multiplication does not happen over all returns is at most $1/n.$ If the posterior for the target edge is multiplied by a factor of at least $(1-\delta)/(c)>(1-\delta)/(1-c)$ each time the algorithm returns to line 1, the bound in the previous paragraph applies for the expected total number of tests in line 1, even if those are conducted with intermittent detours to line 1 in between. If the posterior is not multiplied as above one or more times, the total number of tests is still at most $n.$ Since the probability that the above multiplication does not happen even once is at most $1/n,$ the expected number of tests in line 1 is at most $\frac{1}{\log c^{\prime}}H(X)+\log^{2}n+(1/n)+1$, which is $\frac{1}{\log(1/(1-c))-H(\delta)}H(X)+\Theta(\log^{2}n)$ (since $c^{\prime}=(1-\delta)^{1-\delta}\delta^{\delta}\frac{1}{1-c}$).
Analysis for repetition of each test in line 1 and line 1:
Let a test be repeated $\ell$ times, where $\ell$ is $\Theta(\frac{z}{(1-2\delta)^{2}})$. Let $X=X_{1}+\cdots+X_{\ell}$ where $X_{i}=1$ if the test is not flipped. Then $\nu=\mathbb{E}[X]=(1-\delta)\ell$ and for $\Delta=1-\frac{1}{2(1-\delta)}$, $Pr[X<\ell/2]=Pr[X<(1-\Delta)\nu]\leq e^{-\frac{\nu\Delta^{2}}{2}}$ by Chernoff bound, since $X_{i}$s are $i.i.d.$ Replacing $\nu$ and $\Delta$, we have $e^{-\frac{\nu\Delta^{2}}{2}}=e^{-\frac{(1-\delta)\ell(\frac{1-2\delta}{2(1-\delta)})^{2}}{2}}$, and when $\ell$ is $\Theta(\frac{z}{(1-2\delta)^{2}})$, this probability is less than $e^{-2z}$. Hence with probability at least $1-e^{-2z}$ the majority returns the correct result.
Upper bounding the number of tests in line 1:
Let $z=\log n$ in the previous paragraph. Then, if every test is repeated $\ell$ times, with $\ell$ as $\Theta(\frac{\log n}{(1-2\delta)^{2}})$, with probability at least $1-1/n^{2}$, the majority of the tests returns the correct result. The probability that a noiseless test becomes positive, is less than constant $c$. After repetition the verdict is positive either because the group has at least one infected node, which happens with probability at most $c$, or because the majority verdict is in error, which happens with probability at most $(1/n^{2}).$ By union bound therefore, the majority verdict is positive with probability at most $c+(1/n^{2}).$ The algorithm moves from line 1 to line 1 whenever the majority verdict is positive, and there is always a return to line 1 after such a move. Thus, the probability that the algorithm returns to line 1 at least $(\log n)/(1-c-(1/n^{2}))$ times is
$(c+(1/n^{2}))^{\log n/(1-c-(1/n^{2}))}$ $=(1-(1-c-(1/n^{2}))^{\log n/(1-c-(1/n^{2}))}$
$<(e^{-(1-c-(1/n^{2}))})^{(\log n)/(1-c-(1/n^{2}))}$
$=e^{-(1-c-(1/n^{2}))(\log n)/(1-c-(1/n^{2}))}=1/n.$
Hence, with probability of at least $1-(1/n)$, the total number of tests in this line is upper bounded by any function that is $\Theta(\frac{\log^{2}n}{(1-2\delta)^{2}(1-c-(1/n^{2}))})$. Note that $\Theta(\frac{\log^{2}n}{(1-2\delta)^{2}(1-c-(1/n^{2}))})$ is also $\Theta(\frac{\log^{2}n}{(1-2\delta)^{2}})$ since $c<1/2$. Even when the upper bound does not hold, the number of tests in this line is at most $n.$ Thus, the expected number of test in this line is upper bounded by $n\times(1/n)+$ any function that is $\Theta(\frac{\log^{2}n}{(1-2\delta)^{2}})$ for all large $n$. Thus, the expected number of tests is upper bounded by any function that is $\Theta(\frac{\log^{2}n}{(1-2\delta)^{2}})$ for all large $n$.
We now analyze the posterior every time the algorithm returns to line 1 from line 1. Such return is possible only if the majority outcome is positive. Suppose the majority outcome is positive and for an edge $e$, $e\cap T\neq\phi$ where $T$ is the tested set. Then, $Pr(e=e^{*}\mid T^{r}=-1)=0,$ $Pr(e=e^{*}\mid T^{r}=+1)=\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, where $E^{\prime}$ is the set of edges each of which has at least one node in $T.$ By the algorithm, $w(V\setminus T)\geq 1-c$, so, $\sum_{f\in E^{\prime}}p_{f}\leq c.$ Hence $q(e)\geq(1-\delta)p(e)/c\geq(1-\delta)p(e)/(1-c)$ since $c<1-c.$ Now, if the majority outcome of testing $T$ is positive, and $e^{*}$ is the target edge, either (1) $e^{*}\cap T\neq\phi$ or (2) the majority outcome is erroneous. In the first case, $q(e^{*})\geq(1-\delta)p(e^{*})/(1-c).$ The probability of the second event is at most $1/n^{2}$ for any majority outcome. Since there are at most $n$ tests overall, probability of the second to ever happen is at most $1/n$. Thus, with overall probability at least $1-(1/n)$, every time the algorithm returns to line 1 from line 1, $q(e^{*})\geq(1-\delta)p(e^{*})/(1-c)$.
Upper bounding the number of tests in line 1:
From the statement of the theorem we are proving here, with probability at most $1-O(1/n)$, the target edge has fewer than $u$ nodes. In this line, we repeat each individual test $\Theta(\frac{\log(g(n)u)}{(1-2\delta)^{2}})$ times, where $g(n)$ can be any function such that $\lim_{n\rightarrow\infty}g(n)=\infty.$ Thus, in the analysis for repetition of tests, $z=\log(g(n)u).$ From the analysis for repetition of tests, probability that the majority verdict is erroneous is less than $\frac{1}{2g(n)u}$. Note that $|S|$ nodes are tested in this line, for a set $S$ identified in the previous steps. By Lemma 4.8, $|S|\leq\frac{\tilde{\mu}}{1-2c}$, where $\tilde{\mu}$ is the expected size of the target edge at this stage. If the target edge has $u$ or more nodes, it will still have at most $n$ nodes. Thus, $\tilde{\mu}$ is at most $u$ + $n\times o(1/n)$, which is $O(u).$ Thus, $|S|$ is $O(\frac{u}{1-2c}).$ Thus, 1) the overall number of tests in this line is at most $O(\frac{\log(g(n)u)}{(1-2\delta)^{2}}\frac{u}{1-2c})$, and 2) by union bound, the probability that the majority verdict is erroneous for at least $1$ node in $S$ is $O(1/g(n))$, which implies that with probability at least $1-O(1/g(n))\rightarrow 1$, the algorithm recovers the states of all nodes it tests in this line.
Upper bounding the overall number of tests and the probability that the target edge is not the output:
Combining the upper bounds for the expected number of tests in lines 1, 1 and  1, the expected number of overall tests can be upper bounded by $\frac{1}{\log 1/(1-c)-H(\delta)}H(X)+O(\frac{\log(g(n)u)}{(1-2\delta)^{2}}\frac{\mu}{(1-2c)})+\Theta(\log^{2}n/(1-2\delta)^{2})$ (any functions that are $\Theta(\cdot)$ of the said functions would suffice).
As argued before, line 1 is the only place that the algorithm actually outputs the infectious nodes. The algorithm comes to this line only when the majority verdict of testing $V\setminus S$ is negative in line 1. If the majority verdict is correct, then all positive nodes are in $S$. Once the algorithm reaches line 1, it may not determine the actual set of infected nodes only if either the majority verdict of testing $V\setminus S$ is erroneous the last time the algorithm is in line 1, or the majority verdict is erroneous for one or more nodes of $S$ in line 1. The probability of the first event is at most $1/n^{2}$, and that of the second is at most $O(1/g(n))$. Thus, once the algorithm reaches this line, it may not determine the actual set of infected nodes with probability at most $(1/n^{2})+O(1/g(n))$ by union bound. The above probability is $O(1/\min(n^{2},g(n)).$ The algorithm may also not output any infected node if it exhausts $n$ tests before it reaches or concludes line 1. The probability of this event is $O(1/n)+O(1/g(n))$, which is $O(1/\min(n,g(n)).$ Thus, the probability that the algorithm does not output the target edge is $O(1/\min(n,g(n)).$
The next theorem extends our upper bound for the SNAGT setting to the noisy setting.
@@
Suppose the result of each test is flipped with probability $\delta<1/2$, and for a graph $G$, the probability distribution $\mathcal{D}$ is such that $\Pr_{e\sim\mathcal{D}}(|e|>u)=o(1/n)$ for any constant $\epsilon>0$. Then, there is an algorithm in the SNAGT setting that performs $O(\frac{\log[un]}{(1-2\delta)^{2}}[u\cdot H(X)+u\log n)])$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
Proof Sketch: We employ the same algorithm described in Theorem 7.1, but modify it to address noise by incorporating test repetition and majority voting. Specifically, each test is repeated $O(\frac{\log[dn]}{1-4\delta(1-\delta)})$ times, and the outcome of the test is determined by the majority vote among these repetitions.
By choosing the number of repetitions as above, we ensure that the probability of a single test producing an incorrect majority outcome is at most $\mathrm{err}\ll\frac{1}{un}.$ Applying a union bound over all tests, the probability that any test is incorrect tends to zero as $n$ grows.
Consequently, with high probability, we may assume that all tests are effectively noiseless. Under this assumption, the situation reduces exactly to the noise-free setting of Theorem 7.1. Hence, the same performance guarantees and conclusions from that theorem apply, completing the proof.